Write the equations in matrix form :

x-y=-5

3x+2y=-5


1
Expert's answer
2021-11-02T15:18:09-0400

"\begin{pmatrix}
 1 & -1 \\
 3 & 2
\end{pmatrix}" "\begin{pmatrix}
 x \\
 y
\end{pmatrix}" ="\begin{pmatrix}
 -5 \\
 -5
\end{pmatrix}"